OpenStudy (anonymous):

determine wether the two lines are parallel, perpendicula, or neither y=4x+4 y=3-1/4x

ganeshie8 (ganeshie8):

first, find the slope of each line

OpenStudy (anonymous):

4x nd 3

ganeshie8 (ganeshie8):

slope is just a number

ganeshie8 (ganeshie8):

y = \(\color{red}{4}\)x + 4

ganeshie8 (ganeshie8):

y=3\(\color{red}{-1/4}\)x

ganeshie8 (ganeshie8):

those are the slopes, becoz the given lines are in form : y = \(\color{red}{m}\)x + c where, \(\color{red}{m}\) is the slope

OpenStudy (anonymous):

wich m =3

ganeshie8 (ganeshie8):

no, m is the slope, which is the things attached to 'x'

ganeshie8 (ganeshie8):

not the other one, that we call constant

OpenStudy (anonymous):

-1/4

ganeshie8 (ganeshie8):

thats right, so slope of first line = 4 slope of second line = -1/4

ganeshie8 (ganeshie8):

now, multiply both of them

ganeshie8 (ganeshie8):

what u get ?

OpenStudy (anonymous):

-4

ganeshie8 (ganeshie8):

check again

ganeshie8 (ganeshie8):

product of slopes = \(4 \times \frac{-1}{4}\) = ?

OpenStudy (anonymous):

i do u multiply straight across

ganeshie8 (ganeshie8):

\(\cancel{4} \times \frac{-1}{\cancel{4}} \) = ?

OpenStudy (anonymous):

-1

ganeshie8 (ganeshie8):

thank god ! :)

ganeshie8 (ganeshie8):

so, we got the product of slopes = -1

OpenStudy (anonymous):

so mean

ganeshie8 (ganeshie8):

that means, both lines are perpendicular

OpenStudy (anonymous):

i mean ur mean i know im slow

ganeshie8 (ganeshie8):

here is the rules : 1) lines are perpendicular if 'product of slopes = -1' 2) lines are parallel if slopes are equal

ganeshie8 (ganeshie8):

slow is ok, its good coz u see things very clearly when u go slow

OpenStudy (anonymous):

how do u know if its parallel

ganeshie8 (ganeshie8):

u find the slopes, first if u get both of them equal, then the lines are parallel

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np.. yw :) i knw u dont mean it when u said im so mean hehe ;p

OpenStudy (anonymous):

im serious

ganeshie8 (ganeshie8):

really :'( should i work on fixing it

OpenStudy (anonymous):

its okaay